# PRISM2: Unlocking Multi-Modal General Pathology AI with Clinical Dialogue

### v1

@@ 3 PRISM2
    Robust slide-level representations,
@@
    Zero-shot question answering,
@@
    Interactive diagnostic dialogue, and
@@
    Automatic report generation.
These capabilities informed our data curation strategy, architectural choices, and training methodology, which we detail in the following sections. By targeting these diverse capabilities within a unified framework, we aimed to create a versatile foundation model that addresses multiple clinical workflows in digital pathology.
3.1 Model architecture
The overall architecture of PRISM2 consists of a slide encoder, a language encoder, and a language model, as illustrated in Fig. 2 and described below.
Figure 2: The architecture and training schematic for PRISM2. Sets of whole slide images are summarized as a base embedding (Attention Pooler class token) which is aligned to the representation of the clinical report diagnostic summary (encoded by the language decoder, BioGPT) by minimizing a contrastive loss. The image latent features are also passed to a pretrained large language model (LLM, Phi-3 Mini) along with a prompt from one of four dialogue templates, with the model’s response updated by minimizing the dialogue loss. Training proceeds in two stages: in stage 1, only the slide encoder, attention pooler, language decoder, and LLM adapter weights are updated; in stage 2, only the LLM adapter and LLM weights are updated. For use in downstream tasks, the latent image features produced by the slide encoder are summarized in two embeddings: (1) the features are pooled into a ‘base embedding’, used for contrastive alignment; and (2) the features are refined by the LLM into a ‘diagnostic embedding’, taken from the LLM hidden state at the <|assistant|> token, after inputting the image.
Slide encoder. We adopt the Perceiver [jaegle2021perceiver] architecture for the slide encoder as in PRISM, but with some modifications. Namely, we reduce the number of blocks from 8 to just 1, where a block consists of a cross attention layer between latent queries and tile embeddings followed by 6 self-attention transformer layers for the latent queries. We use Grouped Query Attention (GQA) [ainslie2023gqa] in the tile embedding cross attention. We decrease the number of latents from 512 to 256 and use an attention pooling layer on these to produce the embedding used in the contrastive objective. This embedding is referred to as the base embedding. The Perceiver and attention pooler have 541M and 79M parameters, respectively.
Language encoder. We use BioGPT [luo2022biogpt] as the language encoder but unlike in PRISM, we do not split the network into a unimodal encoder and multi-modal decoder, rather we dedicate the entire network to serve as a language encoder. We initialize a class token which is concatenated with the input text sequence. The final layer hidden state of the class token is then used as the text embedding for the contrastive objective.
Large language model decoder. For the dialogue-focused language model component we use Phi-3 Mini [abdin2024phi] which is a 3.8B parameter decoder-only LLM. We adopt a LLaVa-style [liu2023visual] approach to enabling Phi-3 Mini to process both images and text. The 256 latents outputted by the slide encoder are further processed by a 2-layer multi-layer perceptron (MLP) adapter (29M parameters) to project them into the Phi-3 embedding space and are inserted into the text token embedding sequence as input to the LLM. This allows Phi-3 to integrate the slide and text information via its self-attention layers. Due to the next-token prediction objective of LLMs, we find that the Phi-3 Mini final layer hidden state of the <|assistant|> token encodes a useful slide representation which we refer to as the diagnostic embedding.
3.2 Clinical dialogue
In order to achieve clinically relevant dialogue and reporting capabilities, we generate and curate various formats of text data with GPT-4o [openai2024gpt4ocard] using ground truth clinical diagnostic reports as seed data (see Appendix A.4 for system prompts). This data consists of report rewrites, de-identified patient history, specimen descriptions, and question-answer pairs. We compose these categories of data into various chat-style templates throughout training to simulate dialogue. We create four main templates to use in training:
  1. 1.
    Clinical report generation. Given a task string (e.g. “Write a report”), the model is tasked to generate the summarized clinical report.
  2. 2.
    Question answering (yes / no). The model is tasked to answer a series of binary questions from a user.
  3. 3.
    Question answering (open-ended). The model is tasked to answer a series of open-ended questions from a user.
  4. 4.
@@ 3.2 Clinical dialogue
Specifically, we adopt the Phi-3 [abdin2024phi] chat template using alternating user and assistant tags. The first user message always contains the WSI latents from the slide encoder followed by an initial prompt. In each template, the patient history and specimen description are randomly included, each with a probability of $20\%$, to allow for conditioning the model on this extra context without being overly reliant on it. In total, we create 685K report rewrites for as many specimens and 14M question-answer pairs.
@@
Clinical report rewrites. We use GPT-4o to rewrite clinical diagnostic reports which are subsequently used for training both the language encoder and the language model (see Sec 3.1 for details). Clinical diagnostic reports are a mixture of synoptic reports and free-text notes. We prompt GPT-4o to rewrite all input text, including patient history and specimen description when available, into natural language sentences while retaining all diagnostically relevant findings. These rewrites are typically longer and more informative than the higher level summaries previously used in PRISM [shaikovski2024prism].
@@
Yes-No question-answer pairs. To broaden the training distribution beyond report generation, we use GPT-4o to generate question-answer pairs through a multi-step process. First, we prompt GPT-4o to convert clinical reports into lists of individual findings, similar to MAIRA-2 [bannur2024maira2], which are then filtered to exclude molecular/IHC results. These filtered findings are fed back to GPT-4o with a prompt to generate question-answer pairs for each finding that can be answered with simply “Yes.” or “No.”. We found this process is naturally biased towards creating questions that frequently have the same answer throughout the dataset, as negative findings are typically not mentioned in the reports. For instance, a report positively identifying ductal carcinoma in situ (DCIS) in breast tissue allows GPT-4o to generate a pair such as: “Q: Is DCIS present? A: Yes.” However, a normal breast tissue report does not mention DCIS, so no negative question-answer pair of the form “Q: Is DCIS present? A: No,” is generated, biasing the dataset (and resulting model) toward affirmative responses. Inspired by [goyal2017vqa], to mitigate this bias, we implemented a simple method to mine complementary images for additional question-answer pairs. Specifically, we pair random reports with random questions and prompt GPT-4o to answer based on the provided report. This yields complementary image-question-answer triplets, balancing the data distribution. We include these complementary QA pairs with a sampling rate of $20\%$ during training.
@@
Open-ended question-answer pairs. We also use GPT-4o to generate open-ended question-answer pairs for each finding. We instruct GPT-4o to create the QA pairs such that the questions are not answerable with just yes/no but require more specific answers. Open-ended QA pairs help mimic more realistic model interactions and improve the diversity of the training data.
3.3 Pre-processing of whole slide images
WSIs are processed using a fully-convolutional network to perform tissue segmentation and tiled at 20×20$\times$20 × magnification (0.5 $\mu$m/pixel) with a size of $224\times 224$ pixels, discarding any tiles that are less than 65% tissue. The resulting tiles are then processed by Virchow2 [zimmermann2024virchow2] to produce tile embeddings. We use the 1280-dimensional class token output of the model. During training, up to 100K tiles are loaded per specimen. In cases where specimens have more than 100K tiles, WSIs in the specimen are dropped until the criterion is met. This restriction does not apply during inference.
3.4 Training recipe
@@
Objectives. PRISM2 is trained with both contrastive and autoregressive objectives. We follow [yu2022coca, shaikovski2024prism] for the formulation of the contrastive loss:
$\mathcal{L}_{\text{con}}=-\frac{1}{N}\left(\sum_{i=1}^{N}\log\frac{\exp(\boldsymbol{v}_{i}^{\top}\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\boldsymbol{v}_{i}^{\top}\boldsymbol{t}_{j}/\tau)}+\sum_{i=1}^{N}\log\frac{\exp(\boldsymbol{t}_{i}^{\top}\boldsymbol{v}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\boldsymbol{t}_{i}^{\top}\boldsymbol{v}_{j}/\tau)}\right),$ (1)
@@
where $N$ is batch size, $\tau$ is a learned temperature, and $(\boldsymbol{v}_{i},\boldsymbol{t}_{i})$ are the image and text representations. For the image and text representations we use the base embedding from the slide encoder and the class token embedding from BioGPT after a linear projection and normalization with the $\ell_{2}$ norm. Only the diagnostic report rewrites are used as the text input for the contrastive objective. For the autoregressive objective, we apply the standard cross-entropy loss on next-token predictions from Phi-3 Mini. We use the chat templates as the text input and only apply the loss to the assistant regions of the templates:
@@
Training. We adopt a two stage training approach. In the first stage the slide encoder, attention pooler, LLM adapter, and BioGPT are unfrozen. We freeze Phi-3 Mini during this stage but still keep the autoregressive objective online to help the slide encoder and adapter align to Phi-3. We set $\lambda_{\text{con}}$ to $0.25$ and $\lambda_{\text{chat}}$ to $1.0$. In the second stage, the contrastive objective is discarded and the slide encoder is frozen. The adapter and Phi-3 Mini are unfrozen to allow the language model to better integrate the image input and model the language of pathology. We also reset the optimizer and learning rate schedule for this stage.
@@
We train the model on 56 A100 40GB GPUs with bf16 precision. We use the AdamW [loshchilov2019dadamw] optimizer with a learning rate warmup for 500 steps which is then held constant. We do one epoch of stage 1 training with a post-warmup LR of $1.0\times 10^{-4}$ on all trainable components except the language encoder which uses an LR of $2.0\times 10^{-5}$. We do 16 epochs of stage 2 training. An epoch is defined as one pass over each of the four chat templates per specimen, for all specimens in the training data. Since the contrastive objective aligns the image representation with the report and not with each chat template, each specimen is paired with its report four times in a single epoch.
Data We collected a dataset of 2,350,518 hematoxylin and eosin (H&E) stained whole slide images (WSIs) corresponding to 685,507 specimens, where each specimen is a collection of one or more WSIs with a corresponding clinical report as shown in Fig. 3. The data is collected from Memorial Sloan Kettering Cancer Center (MSKCC), spanning a diverse set of tissues and indications, as well as diverse institutions that submitted cases to MSKCC for review.
Figure 3: PRISM2 training data. a The training data can be grouped on the level of patients, cases, specimens, blocks, or slides, as shown. PRISM2 aligns a specimen-level image representation, consisting of multiple whole slide images, with specimen-level clinical reports. b The top identified tissue types are shown as a fraction of specimens. There are over 150 additional tissue types in the long tail, summarized here as ‘Other Tissues’ (shown gray, along with ‘Unknown Tissue’, on the right). c 36% of all specimens (comprising 14% of all slides) are comprised of slides submitted to MSKCC from diverse institutions across the world. d The most severe diagnosis is shown as a fraction of the training dataset specimens (parsed from clinical reports with ChatGPT). If a clinical report describes the presence of a precursor to cancer along with invasive cancer, then that specimen is labeled cancer here as it is the more severe condition.
3.5 Zero-shot prediction
Both contrastive and generative caption training objectives enable zero-shot prediction for classification tasks, demonstrating the model’s ability to learn clinically relevant tasks directly from diagnostic reports without explicit task-specific training.
Zero-shot prediction with dialogue completion. The question-answering ability of PRISM2 is a straightforward mechanism of zero-shot prediction. Binary classification tasks lend themselves well to yes/no questions (e.g. Is a malignant tumor present?). We refer to this method as yes/no, or ‘YN’, dialogue-based zero-shot prediction. The YN dialogue has the advantage over contrastive zero-shot prediction (see ‘Zero-shot prediction by ranking similarity scores’ below) in that no comprehensive list of negative prompts is required (e.g. keywords covering all non-cancer conditions). Constrained multi-class classification lends itself well to ranking the answers when posing an open-ended question (e.g. What type of carcinoma is suggested by the specimen?). We refer to this method as open-ended or ‘OPN’ dialogue-based zero-shot prediction.
Both YN-, and OPN-based dialogue can quantify the probability of predictions. For each zero-shot classification task, we choose a question prompt $w_{<i}$ (e.g. Is malignant tumor present?) and consider a set $C$ of possible completions $c$. For YN dialogue, these are ‘Yes’ or ‘No’; for OPN dialogue, these may be fill-in-the blank options. For example, when determining a breast cancer subtype, one may complete the phrase “The specimen contains             ” with “invasive ductal carcinoma” or “invasive lobular carcinoma”. Besides the prompt, we provide the model with context $x$, which includes WSI features for the given specimen. When using yes/no questions, the context may also include the diagnostic report generated by PRISM2 (referred to as ‘DYN’ dialogue-based zero-shot), which is appended to the Phi-3 Mini input sequence after the image and before the question. We hypothesize that generating the diagnosis first can help the model to answer a yes/no question. Model predictions are ranked by the normalized log probability of each fill-in-the-blank completion.
Formally, we define our dialogue-based zero-shot classifier as
@@ 3.5 Zero-shot prediction
where $c$ is a completion in a set of completions $C$ that cover all classes in the task, $w_{<i}$ is a tokenized prompt containing a question and $x$ is the context which includes WSI embeddings from the slide encoder and, optionally, a tokenized generated diagnosis. As in  [brown2020gpt3], we normalize log probabilities by the prior (log) probability without context $x$.
Contrastive zero-shot prediction by ranking similarity scores. The contrastive image-text alignment objective can be reformulated as multi-class classification. Given a specimen-level embedding $\boldsymbol{v}$ from the PRISM2 slide encoder and a set of prompt embeddings $\boldsymbol{t}\in\boldsymbol{T}$ from the language encoder, one per each class, we compute the cosine similarity between each prompt embedding and apply a softmax activation to the resulting logits. Each entry in the vector corresponds to a probability of that prompt matching the slide. In the case of multiple per class, we take the prompt with the highest similarity score as the prompt representing that class. Formally, the probability of the class $i$ being the true class is

### v2

@@ 4.2 Creating clinical dialogue for training
    Clinical report generation. Given a task string (i.e. “Write a report”), the model must generate the clinical report.
@@
    Yes-No question answering. Given a binary user question, the model must generate an answer.
@@
    Open-ended question answering. Given an open-ended user question, the model must provide an answer.
@@
    Multiple-choice question answering. Given a user question and a list of answer options, the model must select the correct option.
  5. 5.
@@
We choose these templates to broaden the training distribution beyond the obvious report generation task and enable dialogue-based interactions with the trained model. The process to create the data for each of these tasks is described below and the full system prompts can be found in Supplementary Note S.2.
@@
Clinical report rewrites. Using the synoptic reports and free-text notes, we prompt GPT-4o to rewrite all input text, including patient history and specimen description when available, into natural language sentences while retaining all diagnostically relevant findings.
@@
Yes-No question-answer pairs. Question-answer pairs are generated through a multi-step process. First, GPT-4o is prompted to convert clinical reports into lists of individual findings, similar to MAIRA-2 [5], which are then filtered to exclude molecular and immunohistochemistry (IHC) results. These filtered findings are then used as input to GPT-4o with a prompt to generate question-answer pairs for each finding that can be answered with simply “Yes” or “No”. While effective, we found this process is naturally biased towards creating questions that have the answer “Yes.”. We expect that this is because negative findings are typically not mentioned in the reports. For instance, a report positively identifying DCIS in breast tissue allows GPT-4o to generate a pair such as: “Q: Is DCIS present? A: Yes.” However, a normal breast tissue report does not mention DCIS, so no negative question-answer pair of the form “Q: Is DCIS present? A: No”, is generated, biasing the dataset (and resulting model) toward affirmative responses. Inspired by Goyal et al. [15], to mitigate this bias, we implemented a simple method to mine complementary images for additional questions. Specifically, we pair random reports with random questions and prompt GPT-4o to answer based on the provided report. This yields complementary image-question-answer triplets, balancing the data distribution.
@@
Open-ended question-answer pairs. Open-ended question-answer pairs are also generated from the list of findings described above. We instruct GPT-4o to create the QA pairs such that the questions are not answerable with just yes/no but require more specific answers. Open-ended QA pairs help mimic more realistic model interactions and improve the diversity of the training data.
Multiple-choice question-option-answer triplets. Multiple-choice question-options-answer triplets are generated from the synoptic reports and free-text notes. We prompt GPT-4.1 [26] to first classify a given report into one of 90 possible CAP Cancer Protocol Templates and then to create as many question-options-answer triplets as possible from the report such that they follow the inferred protocol. We only provide the name of the 90 possible protocols and not the full template as we rely on the inherent recall of these protocols by GPT-4.1.
Image-text matching. The matching task uses the clinical report rewrites described above and does not require further preprocessing. This task aims to force the data to use the image information to complete the task.
4.3 PRISM2 architecture
The overall architecture of PRISM2 consists of a slide encoder, a language encoder, and a language model, as illustrated in Extended Data Fig. 1 and described below.
Slide encoder. The slide encoder uses a perceiver [16] architecture. The PRISM2 perceiver uses 1 block, defined as a cross attention layer between latent queries and tile embeddings followed by 6 self-attention transformer layers for the latent queries. acGQA [3] is used in the tile embedding cross attention. The number of latent queries is set to 256 and an attention pooling layer is used on these to produce the embedding used in the contrastive objective. This embedding is referred to as the base embedding. The perceiver and attention pooler have 541M and 79M parameters, respectively.
Language encoder. BioGPT [25] is used as the language encoder. We initialize a learnable class token which is concatenated with the input text sequence. The final layer hidden state of the class token is used as the text embedding for the contrastive objective after a linear projection and $\ell_{2}$ normalization.
Large language model decoder. For the dialogue-focused language model component we use Phi-3 Mini [1], a 3.8B parameter decoder-only LLM. We adopt a LLaVa-style [21] approach to enabling Phi-3 Mini to process both images and text. The 256 latents outputted by the slide encoder are further processed by a 2-layer multi-layer perceptron (MLP) adapter (29M parameters) to project them into the Phi-3 embedding space and are inserted into the text token embedding sequence as input to the LLM. This allows Phi-3 to integrate the slide and text information via its self-attention layers. In addition to decoding the language tokens, we also perform analysis with the Phi-3 Mini final layer hidden state of the <|assistant|> token. We refer to as the diagnostic embedding.
4.4 PRISM2 training
@@
Objectives. PRISM2 is trained with both contrastive and autoregressive objectives. Following [40, 29], the contrastive loss is
$\mathcal{L}_{\text{con}}=-\frac{1}{N}\left(\sum_{i=1}^{N}\log\frac{\exp(\boldsymbol{v}_{i}^{\top}\boldsymbol{t}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\boldsymbol{v}_{i}^{\top}\boldsymbol{t}_{j}/\tau)}\right.+\left.\sum_{i=1}^{N}\log\frac{\exp(\boldsymbol{t}_{i}^{\top}\boldsymbol{v}_{i}/\tau)}{\sum_{j=1}^{N}\exp(\boldsymbol{t}_{i}^{\top}\boldsymbol{v}_{j}/\tau)}\right),$ (1)
@@
where $N$ is batch size, $\tau>0$ is a learned temperature, and $(\boldsymbol{v}_{i},\boldsymbol{t}_{i})$ are the image and text representations, respectively, as defined above. Only the diagnostic report rewrites are used as the text input for the contrastive objective. For the autoregressive objective, we apply the standard cross-entropy loss on next-token predictions from Phi-3 Mini. We use the chat templates as the text input and only apply the loss to the assistant regions of the templates:
@@
Data sampling. For the contrastive objective, report rewrites are randomly sampled from five rewrites per report. For the autoregressive objective, we adopt the Phi-3 [1] chat template using alternating user and assistant tags. The first user message always contains the WSI latents from the slide encoder followed by an initial prompt. In each template, the patient history and specimen description are randomly included, each with a probability of $20\%$, to allow for conditioning the model on this extra context without being overly reliant on it. Complementary QA pairs are included with a sampling rate of $20\%$. In total, we create 685K report rewrites for as many specimens and 14M question-answer pairs.
@@
Optimization. We adopt a two stage training approach. In the first stage the slide encoder, attention pooler, LLM adapter, and BioGPT weights are updated using the full objective. The Phi-3 Mini weights are constant (frozen) during this stage. We set $\lambda_{\text{con}}$ to $0.25$ and $\lambda_{\text{chat}}$ to $1.0$. In the second stage, the contrastive objective is discarded and the slide encoder weights are frozen. The LLM adapter and Phi-3 Mini are updated to allow the language model to better integrate the perceiver latents and model the language of pathology. The optimizer and learning rate schedules are reset for this stage.
@@
We train the model on 56 A100 40GB GPUs with bf16 precision. We use the AdamW [23] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ and a learning rate (LR) warmup for 500 steps which is then held constant. We do one epoch of stage 1 training with a post-warmup LR of $1.0\times 10^{-4}$ on all trainable components except the language encoder which uses an LR of $2.0\times 10^{-5}$, and a 0.1 decay on the slide encoder weights. The large language encoder (Phi-3-Mini) weights are kept frozen in stage 1 but all dialogue objectives are applied. We do 16 epochs of stage 2 training with 500 steps of LR warmup to $2.0\times 10^{-5}$ and no weight decay. The slide and language encoders (perceiver and BioGPT) are frozen and the large language encoder (Phi-3-Mini) is unfrozen for fine-tuning. An epoch is defined as one pass over each of the four chat templates per specimen, for all specimens in the training data. Because the contrastive objective aligns the image representation with the report and not with each chat template, each specimen is paired with its report four times in a single epoch. Several engineering changes were required for efficient training (see Supplementary Note S.3 for details).
Survival fine-tuning. The PRISM2 slide encoder was fine-tuned on a large dataset of 225,597 cases with OS event data, containing 69,114 death events. Survival prediction is trained by learning a new cross-attention pooling in the (perceiver) slide encoder for “survival” embeddings instead of “base” embeddings and adding a linear regression head on top. This head predicts the log relative hazard, along with the event time in months, for the partial log likelihood loss, completing the Cox proportional hazards model. We used the AdamW [23] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.98$ and a learning rate (LR) warmup for 4000 steps which is then held constant. We used a post-warmup LR of $1.0\times 10^{-4}$ for the slide encoder and $1.0\times 10^{-3}$ for the prediction head, with a 0.0001 weight decay. The comparison survival specialist model was trained from scratch using the same architecture with the same hyperparameters as stage 1 PRISM2 training.
4.5 Evaluation methods
Direct prediction with question answering. Direct prediction with question answering refers to a setting without additional training and includes multiple-choice and yes-no settings. With both yes-no and multiple-choice QA, we can quantify the probability of predictions. For each direct classification task, we choose a question prompt $w_{<i}$ (e.g. Is malignant tumor present?) and consider a set $C$ of possible completions $c$. For yes-no QA, these are ‘Yes’ or ‘No’; for multiple-choice QA, these are the letter options (eg. ‘A’, ‘B’).
Formally, we define our dialogue-based direct predictions as
@@ 4.5 Evaluation methods
where $c$ is a completion in a set of completions $C$ that cover all classes in the task, $w_{<i}$ is a tokenized prompt containing a question and $x$ is the context which contains WSI embedding tokens from the slide encoder.
Contrastive direct prediction by ranking similarity scores. We evaluate direct prediction with PRISM and TITAN by reformulating their contrastive image-text alignment objectives as multi-class classification. Given a specimen-level embedding $\boldsymbol{v}$ from the slide encoder and a set of prompt embeddings $\boldsymbol{t}\in\boldsymbol{T}$ from the language encoder, one per each class, we compute the cosine similarity between each prompt embedding and apply a softmax activation to the resulting logits. Each entry in the vector corresponds to a probability of that prompt matching the slide. In the case of multiple per class, we take the prompt with the highest similarity score as the prompt representing that class. Formally, the probability of the class $i$ being the true class is